OpenStudy (anonymous):

the length of a rectangle is six feet more than its width, and the perimeter is 80 feet. find the length and width of the rectangle

OpenStudy (anonymous):

Well first lets write an expression to describe the relationship. Since we are looking for the perimeter of a rectangle, we just take add the length and the width and double. 2L + 2W = 80 It says that the length is six feet more than width, so: L = W + 6 Now we can go ahead and plug in the L back into our original equation 2(W+6) + 2W = 80 2W + 12 + 2W = 80 4W + 12 = 80 4W = 68 W = 17 Remember it said that Length was six more than width? well since we know the width, we can use that to find the length: L = W + 6 L = 17 + 6 L = 23 so your Length is 23 and the Width is 17